A sample of CH4 gas has a volume of 367 mL, has a pressure of 704 mm Hg, and is at a temperature of 7.50 °C. Calculate the amount (in moles) of CH4 in the gas sample. ____mol
Answer #1

We have the relation PV = nRT ………(1)

Here, V = 367 mL = 0.367 L, P = 704 mm Hg = 0.926 atm, T = 7.5 °C = 273 + 7.5 = 280.5 K

R = 0.082 L.atm/K.mol

n = ?

Using relation (1)

n = PV/RT = (0.926 x 0.367)/(0.082 x 280.5) = 0.015 mol

Amount of CH4 in the gas sample = 0.015 mol